OpenStudy (anonymous):

Solve for the variable. a - 16 = -10

jimthompson5910 (jim_thompson5910):

You want to isolate 'a' To do this, you need to move that -16 over to the other side. So you do what to both sides to do this?

OpenStudy (anonymous):

- -16

jimthompson5910 (jim_thompson5910):

good, you can subtract -16 or you can add 16 to both sides

jimthompson5910 (jim_thompson5910):

So a - 16 = -10 a - 16+16 = -10+16 a = 6

OpenStudy (anonymous):

... In order for it to become negative you need (a) to be less than 10.

OpenStudy (anonymous):

@Aurora-Borealis and @jim_thompson5910 thank you guys really helped me out .

jimthompson5910 (jim_thompson5910):

you're welcome